# Picosecond Precision Heavy Ion Detector for hypernuclei lifetimes studies

### v1

@@ Abstract
In this paper, we present the design and preliminary performance evaluation of a new heavy ion detector for direct measurements of heavy hypernuclei lifetimes. The detector employs the previously developed 10 picosecond resolution Radio Frequency (RF) Timer, which converts the temporal information of incident particles into spatial coordinates of secondary or photoelectrons on a position-sensitive detector by means of circular RF scanning in the 500–1000 MHz range. Here, we report the detector design to achieve efficient suppression of accidental background and effective separation of prompt reaction products and delayed events from hypernuclei decays, results of test studies carried out with RF synchronized laser as well as preliminary results obtained by using alpha particles. Dedicated Monte-Carlo simulations have been performed to estimate the detector’s performance under realistic experimental conditions at RF-driven electron, photon, or proton beams. The results confirm the feasibility of the proposed design and provide a basis for upcoming experimental measurements, based on the delayed fission detection.
@@ 1 Introduction
A more advanced experiment at KEK later employed the (, ) reaction, coupled with -weak decay [8], to positively identify hypernuclei from their ground and low-lying excited states, particularly in heavier targets. Using the time-of-flight (TOF) technique and a production reference time, the experiment directly measured decay time spectra for B, C, Al, Si, and Fe. This method, combining precise hypernuclear identification with direct decay time measurement, yielded the most reliable results, showing lifetimes of ps, ps, ps, ps and ps respectively [9]. These findings suggest a saturation of hypernuclear lifetime for . However, extending this approach to even heavier hypernuclei faces significant challenges due to the dramatically lower production yields for low-lying hypernuclei as A increases further. Heavy hypernuclear lifetimes measurements using the delayed fission technique, combined with the recoil shadow method, were first performed in Kharkov [10, 11], where hypernuclei were produced via electron bombardment of bismuth targets, and at CERN through antiproton interactions with bismuth and uranium targets [12]. Reported lifetimes for hypernuclei in the bismuth region were ns [10], ps [11] and ps [12], while those for uranium hypernuclei were significantly shorter, at ps [12]. To clarify these conflicting results, the COSY-13 collaboration conducted subsequent experiments, using a similar method and employing a 1.5 GeV proton beam on gold, bismuth, and uranium targets [13, 14, 15]. The goal was to increase the momentum transfer to the hypernuclei, thereby enhancing their decay range and reducing measurement uncertainty. The reported fitted lifetime was ps, averaged Au, Bi and U. However, these lifetime measurements were indirect, relying heavily on theoretical models to describe the system’s time evolution.
Another lifetime experiment was conducted at JLab [16], providing a direct measurement of the lifetimes of medium-heavy hypernuclei. These hyper-fragments were produced via fission or breakup of heavy hypernuclei initially generated by a 2.34 GeV photon beam incident on thin Fe, Cu, Ag, and Bi target foils. For each event, coincident fragment pairs were detected using a low-pressure multi-wire proportional chamber system [17]. The lifetime was determined from the decay time spectrum, calculated as the time difference between the zero points of the detected fragment pairs. The measured lifetime for each target represents a statistical average over a range of hypernuclear masses, with the mean mass being approximately half that of the target material. The results indicate a lifetime of about 200 ps, with statistical errors ranging from 3.5 ps to 49.3 ps and a systematic error of 10 ps.
@@
We present the design and preliminary performance evaluation of a new heavy ion detector for direct measurements of heavy hypernuclei lifetimes at RF-driven electron, photon or proton beams. The system integrates the Radio Frequency Timer (RFT) as its core timing device and has been optimized to suppress accidental background and enable efficient separation of prompt and delayed events. This work focuses on the detector system. The hypernuclear experiments themselves will be discussed in a separate publication. We describe the detector concept in section 2, the laboratory testing in section 3, and Monte Carlo simulations used to estimate performance under realistic beam conditions in section 4, followed by a summary and outlook in section 5.
@@ 2 Heavy Ion Detector Concept and Design
However, among the nuclear reaction products are also hypernuclei, which are stopped within the target material. These hypernuclei undergo weak decays after characteristic lifetimes of approximately 200 ps, emitting charged particles (pions, protons) that subsequently cross the target surface and produce delayed SEs. These delayed SEs are the key signal for hypernuclear lifetime measurements. Accurate reconstruction of the hypernuclei lifetime requires precise measurement of the time distribution of these delayed SEs.
The main challenge lies in distinguishing delayed SEs associated with hypernuclear decay from the overwhelming background of prompt SEs generated directly by the primary beam or by prompt nuclear reactions. In the case of heavy hypernuclei, this separation is significantly simplified: the decay of a hyperon within such a nucleus induces nuclear fission, producing two fission fragments emitted nearly back-to-back (at 180∘). Each of these fragments generates 100 SEs upon crossing the target surface. This SE yield is roughly two orders of magnitude higher than that of a typical primary beam particle. The ratio of delayed and instantaneous secondary electrons associated with prompt fission of nuclei induced by the primary particles will be determined by the probability of the formation of hypernuclei in nuclear reactions. In the case of interactions of 3 GeV pions and protons with an emulsion, this is about 0.09 percent [22] and 2 percent in the case of interactions of 1.5 GeV photons with Bi [23].
@@ 3 Laboratory Tests
Notably, measurements with a monolayer graphene photoemitter revealed a time distribution closely resembling the expected lifetime spectrum of heavy hypernuclei decays. A significant number of prompt photoelectrons were detected, while smaller part of the events exhibited delays of up to 2 nanoseconds. Fitting the delayed component with an exponential function yielded a lifetime of ps (figure 5 bottom), comparable to theoretical predictions for hypernuclear lifetime saturation. This result demonstrates the detector’s capability to resolve lifetime spectra with the desired precision.
@@ 4 Expected lifetime measurement performance: Monte-Carlo simulations
The ratio of prompt and delayed fission events is estimated to be [23]. Besides that, the secondary electrons generated by the initial beam particles will also be registered as prompt events, so total ratio can vary from to depending on the experimental conditions, specifically the type and properties of the initial beam. The measured prompt events times will follow a Gaussian distribution around the time zero point with a standard deviation determined by the detector’s time resolution, while the delayed events will follow an exponential distribution superimposed with Gaussian noise (again with standard deviation determined by the detector’s time resolution).
@@ 5 Summary
A new time-resolved heavy ion detector, based on circular radio-frequency scanning of secondary electrons, has been designed, built and tested in the laboratory using a synchronized femtosecond laser. The measured time resolution of the detector is approximately 12 ps. This system is particularly suitable for hypernuclear lifetime studies, enabling the separation of high-rate prompt reaction products from delayed events of interest. Laboratory studies of hot carrier lifetimes in graphene, which exhibit a time structure similar to that expected for hypernuclear decays, demonstrate the detector’s capability for such measurements. Additionally, the detection of secondary electrons generated by alpha particles has been successfully verified. Detailed Monte Carlo simulations confirm that the expected statistical uncertainty for direct hypernuclear lifetime measurements with this detector is about 10 ps under realistic conditions. These results indicate that the proposed detector concept is feasible and provides a solid basis for future experiments with electron, photon, or proton beams.

### v2

# Picosecond Precision Heavy Ion Detector for Hypernuclei Lifetime Studies

@@ Abstract
In this paper, we present the design and preliminary performance evaluation of a new heavy ion detector for direct measurements of heavy hypernuclei lifetime. The detector employs the previously developed 10 picosecond resolution Radio Frequency (RF) Timer, which converts the temporal information of incident particles into spatial coordinates of secondary or photoelectrons on a position-sensitive detector by means of circular RF scanning in the 500–1000 MHz range. Here, we report the detector design to achieve efficient suppression of accidental background and effective separation of prompt reaction products and delayed events from hypernuclei decays, results of test studies carried out with RF synchronized laser as well as preliminary results obtained by using alpha particles. Dedicated Monte-Carlo simulations have been performed to estimate the detector’s performance under realistic experimental conditions at RF-driven electron, photon, or proton beams. The results confirm the feasibility of the proposed design and provide a basis for upcoming experimental measurements, based on the delayed fission detection.
@@ 1 Introduction
A more advanced experiment at KEK later employed the (, ) reaction, coupled with -weak decay [8], to positively identify hypernuclei from their ground and low-lying excited states, particularly in heavier targets. Using the time-of-flight (TOF) technique and a production reference time, the experiment directly measured decay time spectra for B, C, Al, Si, and Fe. This method, combining precise hypernuclear identification with direct decay time measurement, yielded the most reliable results, showing lifetimes of ps, ps, ps, ps and ps respectively [9]. These findings suggest a saturation of hypernuclear lifetime for . However, extending this approach to even heavier hypernuclei faces significant challenges due to the dramatically lower production yields for low-lying hypernuclei as A increases further. Heavy hypernuclear lifetime measurements using the delayed fission technique, combined with the recoil shadow method, were first performed in Kharkov [10, 11], where hypernuclei were produced via electron bombardment of bismuth targets, and at CERN through antiproton interactions with bismuth and uranium targets [12]. Reported lifetimes for hypernuclei in the bismuth region were ns [10], ps [11] and ps [12], while those for uranium hypernuclei were significantly shorter, at ps [12]. To clarify these conflicting results, the COSY-13 collaboration conducted subsequent experiments, using a similar method and employing a 1.5 GeV proton beam on gold, bismuth, and uranium targets [13, 14, 15]. The goal was to increase the momentum transfer to the hypernuclei, thereby enhancing their decay range and reducing measurement uncertainty. The reported fitted lifetime was ps, averaged Au, Bi and U. However, these lifetime measurements were indirect, relying heavily on theoretical models to describe the system’s time evolution.
Another lifetime experiment was conducted at JLab [16], providing a direct measurement of the lifetime of medium-heavy hypernuclei. These hyper-fragments were produced via fission or breakup of heavy hypernuclei initially generated by a 2.34 GeV photon beam incident on thin Fe, Cu, Ag, and Bi target foils. For each event, coincident fragment pairs were detected using a low-pressure multi-wire proportional chamber system [17]. The lifetime was determined from the decay time spectrum, calculated as the time difference between the zero points of the detected fragment pairs. The measured lifetime for each target represents a statistical average over a range of hypernuclear masses, with the mean mass being approximately half that of the target material. The results indicate a lifetime of about 200 ps, with statistical errors ranging from 3.5 ps to 49.3 ps and a systematic error of 10 ps.
@@
We present the design and preliminary performance evaluation of a new heavy ion detector for direct measurements of heavy hypernuclei lifetime at RF-driven electron, photon or proton beams. The system integrates the Radio Frequency Timer (RFT) as its core timing device and has been optimized to suppress accidental background and enable efficient separation of prompt and delayed events.
@@ 2 Heavy Ion Detector Concept and Design
However, among the nuclear reaction products are also hypernuclei, which are stopped within the target material. These hypernuclei undergo weak decays after characteristic lifetime of approximately 200 ps, emitting charged particles (pions, protons) that subsequently cross the target surface and produce delayed SEs. These delayed SEs are the key signal for hypernuclear lifetime measurements. Accurate reconstruction of the hypernuclei lifetime requires precise measurement of the time distribution of these delayed SEs.
The main challenge lies in distinguishing delayed SEs associated with hypernuclear decay from the overwhelming background of prompt SEs generated directly by the primary beam or by prompt nuclear reactions. In the case of heavy hypernuclei, this separation is significantly simplified: the decay of a hyperon within such a nucleus induces nuclear fission, producing two fission fragments emitted nearly back-to-back (at 180∘). Each of these fragments generates 100 SEs upon crossing the target surface. This SE yield is roughly two orders of magnitude higher than that of a typical primary beam particle. The ratio of delayed and instantaneous secondary electrons associated with prompt fission of nuclei induced by the primary particles will be determined by the probability of the formation of hypernuclei in nuclear reactions. In the case of interactions of 3 GeV pions and protons with an emulsion, this is about 0.09 percent [22] and 0.5 percent in the case of interactions of 1.5 GeV electrons with Bi [23] (see more details in 4).
@@ 3 Laboratory Tests
Notably, measurements with a monolayer graphene photoemitter revealed a time distribution closely resembling the expected lifetime spectrum of heavy hypernuclei decays. A significant number of prompt photoelectrons were detected, while smaller part of the events exhibited delays of up to 2 nanoseconds. These extended photoemission times in graphene can be theoretically attributed to its high optical phonon energy and limited phase space for acoustic phonon scattering, which result in hot carriers lifetimes ranging from hundreds of picoseconds to a few nanoseconds [26, 27]. Experimentally, the nanosecond-scale photoemission times in graphene at room temperature were first observed using a highly sensitive time-of-flight angle-resolved photoemission spectroscopy (TOF ARPES) technique operated at very low laser fluence (10 J/cm2) [27].
Although in our studies the mean value of the delayed photoemission component varied across graphene samples, an exponential tail was consistently observed. For one of the samples, the result of the fit corresponds to a lifetime of about ps (figure 5 bottom), comparable to theoretical predictions for hypernuclear lifetime saturation. We fit the prompt part of the graphene spectrum with the Gaussian function and then fit the tail, starting 3 sigma from the Gaussian mean, with an exponential function. The parameter values shown in figure 5 (bottom) correspond to the exponential fit. This result demonstrates the detector’s capability to resolve lifetime spectra with the desired precision.
@@
4 Fission Yields and Secondary-Electron Production in the Context of -hypernuclear Lifetime Experiment Feasibility
For the estimation of prompt and delayed event rates, we considered the following configuration: 1.5 GeV electrons delivered in 500 MHz bunches with a duration of 2 ps irradiate a thin bismuth foil of thickness 4 mg/cm2, corresponding to T=0.00064 radiation lengths. At these beam energies and for such a thin target, electron–nucleus interactions are dominated by virtual-photon processes. Electron straggling due to bremsstrahlung can therefore be neglected. The yield Y of a photoproduction process with cross section (k), induced by a single electron through a virtual photon of energy k, is
where N is Avogadro’s number, A=209 is the atomic number of bismuth, X0 = 6.3 g/cm2 is the radiation length, and teq=0.017 is the radiation length of the equivalent radiator, which represents the virtual photon flux of the incident electron [28].
In the 1.5 GeV energy region, lifetime measurements of -hypernuclei can be performed using a high-energy, high-repetition, photon-induced reactions for which prompt fission cross section is approximately 2.5 mb. Delayed fission is associated with the production of a -hypernuclei and its subsequent decay, occurring on a timescale of about 200 ps. The dominant production channels are
and, considering only the first reaction, the delayed-fission cross section for bismuth is roughly 12.5 b (see [23] and references therein).
Assuming prompt and delayed photo-fission cross sections of 2.5 mb and 12.5 b, respectively, in the 1.0–1.5 GeV energy range, a flux of 1012 electrons traversing a 4 mg/cm2 Bi target is expected to produce, on average, one delayed-fission event and roughly 200 prompt-fission events. Each fission event emits two oppositely directed fragments which, upon exiting the target, generate more than 100 secondary electrons per fragment (360 according to [23]). In contrast, a 1.5 GeV beam electron produces only 0.03 secondary electrons in the forward direction and 0.015 in the backward direction [29].
For a beam intensity of 1010 e/s, each 2 ns period delivers a 2 ps bunch containing 20 electrons, generating approximately 0.6 forward and 0.3 backward secondary electrons. A beam of this intensity yields about 200 prompt-fission events and at least one delayed-fission event within 100 s, with each fission event producing more than 100 secondary electrons on each side of the target. The time distribution of these secondary electrons is recorded with the HID system. Additionally, by using the PSD design described in section 2 (see figure 2), the detector can be made effectively insensitive to prompt events, allowing only the low-rate delayed candidate hypernuclear-production events to be registered and processed by the data acquisition system.
Under these conditions, the secondary-electron yield from fission fragments exceeds that from primary beam electrons by more than two orders of magnitude. Therefore, efficient suppression of events induced by beam particles can be achieved through appropriate collimation of the secondary electrons and suitable operating conditions of the detection system. Over 100 hours of beam time, over 3000 delayed-fission events can be accumulated, allowing a determination of the -hypernuclei lifetime with a statistical precision under 3%.
@@ 5 Expected lifetime measurement performance: Monte-Carlo simulations
As discussed in section 4, the ratio of prompt to delayed fission events is estimated to be 200:1 for a bismuth target, while prompt secondary electrons generated directly by the primary beam can be suppressed through collimation in the detector system. It was also shown that under certain conditions, approximately 3000 delayed-fission events can be collected over 100 hours of beam time. In general, the number of detected delayed-fission events () and the prompt-to-delayed ratio () will vary depending on the specific experimental conditions: the type and properties of the initial beam, the target and the actual beam time.
For the simulation estimates presented below, we assume and prompt-to-delayed ratios 103 and 105. The measured prompt events times will follow a Gaussian distribution around the time zero point with a standard deviation determined by the detector’s time resolution, while the delayed events will follow an exponential distribution superimposed with Gaussian noise (again with standard deviation determined by the detector’s time resolution).
@@ 6 Summary
A new time-resolved heavy ion detector, based on circular radio-frequency scanning of secondary electrons, has been designed, built and tested in the laboratory using a synchronized femtosecond laser. The measured time resolution of the detector is approximately 12 ps. This system is particularly suitable for hypernuclear lifetime studies, enabling the separation of high-rate prompt reaction products from delayed events of interest. Laboratory studies of hot carrier lifetime in graphene, which exhibit a time structure similar to that expected for hypernuclear decays, demonstrate the detector’s capability for such measurements. Additionally, the detection of secondary electrons generated by alpha particles has been successfully verified. Detailed Monte Carlo simulations confirm that the expected statistical uncertainty for direct hypernuclear lifetime measurements with this detector is about 10 ps under realistic conditions. These results indicate that the proposed detector concept is feasible and provides a solid basis for future experiments with electron, photon, or proton beams.